# Self-Normalized Resets for Plasticity in Continual Learning

### v2

@@ Appendix C Additional Experimental Details and Hyperparameter Sweep
Let $T$ be the $1-\frac{\lambda}{p-\lambda}$ percentile of a Geometric$(p)$ distribution. Then the optimal hypothesis test takes the form $X_{\tau}=\mathbf{1}\{Z_{\tau}=0\}$ where $\tau={\rm min}(s:Z_{s}=1)\wedge T$.
@@
Let $F$ be the CDF of the Geometric$(p)$ distribution. Let $T^{*}$ be the $1-\frac{\lambda}{p-\lambda}$ percentile of the Geometric$(p)$ distribution. Note, since $\lambda<\frac{p}{2}$ then $1-\frac{\lambda}{p-\lambda}\in(0,1)$ and is a valid percentile. Then for any $T\geq T^{*}$ we have that
@@
For notational convenience, define $\bar{\alpha}=\alpha(1-\alpha)^{-1}$. Notice that by Proposition 2.1, under SNR, neuron $i$, $(I=1,2)$, is reset if it inactive for any longer that time $\bar{T}=\log(\bar{\alpha}))/\log(1-p_{i})$. Consequently, the expected delay penalty, $\mathbb{E}[\tau|H_{0}]+\mathbb{E}[\tau|H_{1}]$ for neuron $i$, is simply

### v3

@@ D.7 Negative Results
Let $T$ be the $1-\frac{\lambda}{p-\lambda}$ percentile of a Geometric(p)(p) distribution. Then the optimal hypothesis test takes the form $X_{\tau}=\mathbf{1}\{Z_{\tau}=0\}$ where $\tau={\rm min}(s:Z_{s}=1)\wedge T$.
@@
Let $F$ be the CDF of the Geometric(p)(p) distribution. Let $T^{*}$ be the $1-\frac{\lambda}{p-\lambda}$ percentile of the Geometric(p)(p) distribution. Note, since $\lambda<\frac{p}{2}$ then $1-\frac{\lambda}{p-\lambda}\in(0,1)$ and is a valid percentile. Then for any $T\geq T^{*}$ we have that
@@
For notational convenience, define $\bar{\alpha}=\alpha(1-\alpha)^{-1}$. Notice that by Proposition D.4, under SNR, neuron $i$, $(I=1,2)$, is reset if it inactive for any longer that time $\bar{T}=\log(\bar{\alpha}))/\log(1-p_{i})$. Consequently, the expected delay penalty, $\mathbb{E}[\tau|H_{0}]+\mathbb{E}[\tau|H_{1}]$ for neuron $i$, is simply